# Simple homotopy types of even dimensional manifolds

### v2

@@ 6.2. The Whitehead torsion
By Lemma 6.3 there is a homotopy automorphism $\bar{f}_{2}\colon L\rightarrow L$ such that $f\circ i_{2}\simeq i_{2}\circ\bar{f}_{2}$. Let $\overline{h}\colon L\rightarrow L$ be the homotopy inverse of $\bar{f}_{2}$, and let $h:=\operatorname{Id}_{S^{1}}\times\overline{h}\colon S^{1}\times L\rightarrow S^{1}\times L$. Note that $h\circ i_{2}=i_{2}\circ\overline{h}\colon L\to S^{1}\times L$. By Corollary 2.33 $\tau(h)=0$. Let $f^{\prime}=f\circ h$. Then by Proposition 2.31, we have that $\tau(f^{\prime})=\tau(f)$. Moreover, combining the above observations we have
$f^{\prime}\circ i_{2}=f\circ h\circ i_{2}=f\circ i_{2}\circ\overline{h}\simeq i_{2}\circ\bar{f}_{2}\circ\overline{h}\simeq i_{2}.$
This means that after changing $f^{\prime}$ by a homotopy we can assume that $f^{\prime}\big{|}_{L}=\operatorname{Id}$.
The embedding $i_{1}\colon S^{1}\rightarrow S^{1}\times L$ represents $(1,0)\in\pi_{1}(S^{1}\times L)\cong C_{\infty}\times C_{m}$. Since $\pi_{1}(f^{\prime})$ is an automorphism of $C_{\infty}\times C_{m}$, $\pi_{1}(f^{\prime})^{-1}(1,0)=(a,b)$ for some $a\in\left\{\pm 1\right\}$, $b\in C_{m}$. It follows from Lemma 6.2 that there is a diffeomorphism $h\in\operatorname{Diff}(S^{1}\times L)$ such that $\pi_{1}(h)(1,0)=(a,b)$. Let $f^{\prime\prime}=f^{\prime}\circ h$. Then $\tau(f^{\prime\prime})=\tau(f^{\prime})$ by Proposition 2.31 and Chapman’s Theorem 2.6. Moreover $\pi_{1}(f^{\prime\prime})(1,0)=(1,0)$, i.e. $f^{\prime\prime}\circ i_{1}\simeq i_{1}$. The diffeomorphisms constructed in Lemma 6.2 keep $L$ pointwise fixed, so $f^{\prime\prime}\big{|}_{L}=\operatorname{Id}$, and after applying a homotopy we can also assume $f^{\prime\prime}\big{|}_{S^{1}}=\operatorname{Id}$.
Finally, we recursively construct maps $f^{\prime\prime}_{j}$ for every $21\leq j\leq 2k-$ 2 such that $f^{\prime\prime}\simeq f^{\prime\prime}_{j}$ rel $S^{1}\vee L$ and $f^{\prime\prime}_{j}\big{|}_{L\cup\operatorname{sk}_{j}(L\times S^{1})}=\operatorname{Id}$. We start with $f^{\prime\prime}_{1}=f^{\prime\prime}$. If $f^{\prime\prime}_{j-1}$ is already defined, we consider the single $j$-cell $E^{\prime}_{j}$ in $S^{1}\times L-L$. Since $f^{\prime\prime}_{j-1}\big{|}_{\partial E^{\prime}_{j}}=\operatorname{Id}$, the cell $E^{\prime}_{j}$ and $f^{\prime\prime}_{j-1}(E^{\prime}_{j})$ together form a map $S^{j}\rightarrow S^{1}\times L$. This map is nullhomotopic (because $\pi_{j}(S^{1}\times L)\cong\pi_{j}(\mathbb{R}\times S^{2k-1})=0$ for $22\leq j\leq 2k-$ 2), so $f^{\prime\prime}_{j-1}\big{|}_{E^{\prime}_{j}}$ is homotopic to $\operatorname{Id}_{E^{\prime}_{j}}$ rel $\partial E^{\prime}_{j}$. We can extend this homotopy to a homotopy between $f^{\prime\prime}_{j-1}$ and an $f^{\prime\prime}_{j}$ with $f^{\prime\prime}_{j}\big{|}_{L\cup\operatorname{sk}_{j}(L\times S^{1})}=\operatorname{Id}$. Therefore $f^{\prime\prime}_{j}$ can be defined for every $21\leq j\leq 2k-$ 2.
To complete the proof of Lemma 6.9, we take $g=f^{\prime\prime}_{2k-2}\colon S^{1}\times L\to S^{1}\times L$. ∎
Proof of Theorem 6.6.
Consider a homotopy equivalence $f\colon S^{1}\times L\to S^{1}\times L$. By Lemma 6.9, we can replace $f$ by a homotopy equivalence $g\colon S^{1}\times L\to S^{1}\times L$ such that $\tau(f)=\tau(g)$ and $g\big{|}_{L\cup\operatorname{sk}_{2k-2}(L\times S^{1})}=\operatorname{Id}$. Then $g$ satisfies the hypotheses of Lemma 6.8, and so $\tau(g)=0$. Therefore $\tau(f)=0$, and $f$ is a simple homotopy equivalence, as desired. ∎
7. The proof of Theorems B, C, LABEL:, and D
We can now combine the results of the previous sections to prove Theorems B, C, LABEL:, and D. Let $n=2k\geq 4$ be an even integer and fix a category $\operatorname{CAT}$ satisfying 1.1.
Recall the definition of the sets $\mathcal{M}^{h}_{s}(M)$, $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$ and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$ from Definition 4.10. Also recall the definition of $A_{2k}(m)$ (Definition 6.4): $A_{2k}(m)\leq\operatorname{Aut}(C_{\infty}\times C_{m})$ denotes the subgroup of matrices $\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$ such that $c^{k}\equiv\pm 1$ mod $m$. Note that $A_{2k}(m)$ acts on $\operatorname{Wh}(C_{\infty}\times C_{m})$ and on the subgroups $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\mathcal{J}_{n}(C_{\infty}\times C_{m})$.
@@ 7. The proof of Theorems B, C, LABEL:, and D
Let $m\geq 2$ and $q_{1},\ldots,q_{k}$ such that $\gcd(m,q_{j})=1$. Let $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ be an $(n-1)$-dimensional lens space with $\pi_{1}(L)\cong C_{m}$. Let $G:=C_{\infty}\times C_{m}\cong\pi_{1}(S^{1}\times L)$. Then there are bijections of pointed sets $\mathcal{M}^{h}_{s}(S^{1}\times L)\cong\mathcal{J}_{n}(G)/A_{2k}(m)$, $\mathcal{M}^{\operatorname{hCob}}_{s}(S^{1}\times L)\cong\mathcal{I}_{n}(G)/A_{2k}(m)$ and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(S^{1}\times L)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/A_{2k}(m)$.
@@
Firstly note that, since $C_{\infty}\times C_{m}$ is polycyclic, it is a good group [FQ, Freedman-book-goodgroups], so 1.1 is satisfied. Moreover, since the map $\psi$ is surjective for $G=C_{\infty}\times C_{m}$, $w\equiv 1$ and $n$ even by Proposition 3.12, the vertical maps in the diagram of Theorem 4.11 (c) are bijections if $M=S^{1}\times L$. Therefore $\mathcal{M}^{h}_{s}(S^{1}\times L)\cong\mathcal{J}_{n}(G)/\operatorname{hAut}(S^{1}\times L)$, $\mathcal{M}^{\operatorname{hCob}}_{s}(S^{1}\times L)\cong q(\mathcal{I}_{n}(G))$, and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(S^{1}\times L)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/\operatorname{hAut}(S^{1}\times L)$.
@@
By Theorem 6.6, every homotopy automorphism of $S^{1}\times L$ is simple, therefore the action of $\operatorname{hAut}(S^{1}\times L)$ on $\operatorname{Wh}(G)$ factors through the action of $\operatorname{Aut}(G)$ (see Definition 4.3 and Remark 4.18). By Theorem 6.5, the image of $\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G)$ is $A_{2k}(m)$. In particular, the orbits of the action of $\operatorname{hAut}(S^{1}\times L)$ and $A_{2k}(m)$ on $\mathcal{J}_{n}(G)$, $\mathcal{I}_{n}(G)$, and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))$ coincide. This implies that $\mathcal{J}_{n}(G)/\operatorname{hAut}(S^{1}\times L)=\mathcal{J}_{n}(G)/A_{2k}(m)$, $q(\mathcal{I}_{n}(G))=\mathcal{I}_{n}(G)/A_{2k}(m)$. Hence, as required, we have $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/\operatorname{hAut}(S^{1}\times L)=\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/A_{2k}(m)$. ∎
@@
In particular Proposition 7.1 implies that $|\mathcal{M}^{h}_{s}(S^{1}\times L)|$, $|\mathcal{M}^{\operatorname{hCob}}_{s}(S^{1}\times L)|$ and $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(S^{1}\times L)|$ are independent of the choice of the $q_{j}$ and of $\operatorname{CAT}$, and only depend on $n$ and $m$. This proves part (a) of Theorems B, C, LABEL:, and D.
@@
From now on, we will write $M^{n}_{m}$ for $S^{1}\times L$, where $L$ is any $(n-1)$-dimensional lens space with $\pi_{1}(L)\cong C_{m}$.
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=1$ if and only if $\mathcal{J}_{n}(G)=0$.
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=\infty$ if and only if $|\mathcal{J}_{n}(G)|=\infty$.
@@
    If $\mathcal{M}^{h}_{s}(M^{n}_{m})$ is finite, then $\frac{|\mathcal{J}_{n}(G)|}{2m^{2}}<|\mathcal{M}^{h}_{s}(M^{n}_{m})|\leq|\mathcal{J}_{n}(G)|$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|=1$ if and only if $\mathcal{I}_{n}(G)=0$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|=\infty$ if and only if $|\mathcal{I}_{n}(G)|=\infty$.
@@
    If $\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})$ is finite, then $\frac{|\mathcal{I}_{n}(G)|}{2m^{2}}<|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|\leq|\mathcal{I}_{n}(G)|$.
@@
    $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|=1$ if and only if $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))=0$.
@@
    $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|=\infty$ if and only if $|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))|=\infty$.
@@
    If $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})$ is finite, then $\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))|}{2m^{2}}<|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|\leq|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))|$.
@@
(a) By Proposition 7.1, $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=|\mathcal{J}_{n}(G)/A_{2k}(m)|$. The group $A_{2k}(m)\leq\operatorname{Aut}(G)$ acts on $\operatorname{Wh}(G)$, and hence on $\mathcal{J}_{n}(G)$, by automorphisms. So $0\in\mathcal{J}_{n}(G)$ is a fixed point of this action, and $|\mathcal{J}_{n}(G)/A_{2k}(m)|=1$ if and only if $\mathcal{J}_{n}(G)=0$.
@@
The proofs of parts (d), (e), and (f) (resp. parts (g), (h) and (i)) are entirely analogous to those of parts (a), (b), and (c) respectively, and so will be omitted for brevity. ∎
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$.
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free.
@@
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|\to\infty$ as $m\to\infty$ (uniformly in $n$).
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|\to\infty$ as $m\to\infty$ (uniformly in $n$).
@@
    There are infinitely many $m$ such that $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|=1$ for every $n$.
@@
    $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|$ is finite for every $n$ and $m$.
@@
    $\limsup_{m\to\infty}\Bigl{(}\inf_{n}|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|\Bigr{)}=\infty$.

### v3

@@ 6. Homotopy automorphisms of $S^{1}\times L$
By Lemma 6.1 the subgroup of realisable automorphisms is generated by $\pi_{1}(f\times g)$ and $\pi_{1}(r_{h})$ for all $f\in\operatorname{hAut}(S^{1})$, $g\in\operatorname{hAut}(L)$ and $h\colon S^{1}\rightarrow\operatorname{Map}(L)$. By our earlier observations and Lemma 6.3, the automorphisms realised as $\pi_{1}(f\times g)$ are precisely those of the form $\left(\begin{smallmatrix}a&0\\ 0&c\end{smallmatrix}\right)$ with $a\in\{\pm 1\}$ and $c^{k}\equiv\pm 1$ mod $m$, and $\pi_{1}(r_{h})$ is of the form $\left(\begin{smallmatrix}1&b\\ 0&1\end{smallmatrix}\right)$, showing that $\operatorname{Im}(\pi_{1})\leq A_{2k}(m)$. On the other hand, by Lemma 6.4 the matrix $\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)$ is realisable, and together with the matrices $\left(\begin{smallmatrix}a&0\\ 0&c\end{smallmatrix}\right)$ with $a\in\{\pm 1\}$ and $c^{k}\equiv\pm 1$ mod $m$ they generate $A_{2k}(m)$, therefore $\operatorname{Im}(\pi_{1})\geq A_{2k}(m)$. ∎
7. The proof of Theorems A, C, D, LABEL:, and E
We can now combine the results of the previous sections to prove Theorems A, C, D, LABEL:, and E. Let $n=2k\geq 4$ be an even integer and fix a category $\operatorname{CAT}$ satisfying 1.1.
We begin by establishing the following key fact concerning the map $\psi$ from the exact sequence (3.2) for the groups $C_{\infty}\times C_{m}$.
@@ 7. The proof of Theorems A, C, D, LABEL:, and E
Let $m\geq 2$. For any even integer $n=2k$, the map
$\psi\colon L_{n+1}^{\mathrm{h}}(\mathbb{Z}[C_{\infty}\times C_{m}])\to\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$
is surjective.
@@
Let $G_{m}:=C_{\infty}\times C_{m}$. We verify that the forgetful map $F\colon L^{\mathrm{s}}_{2k}(\mathbb{Z}G_{m})\to L^{\mathrm{h}}_{2k}(\mathbb{Z}G_{m})$ is injective. The conclusion then follows from the exact sequence (3.2). First we apply Shaneson splitting to the domain and the codomain, to obtain a commutative diagram whose rows are split short exact sequences and whose vertical maps are the forgetful maps [Ra86].
[…]
By the five lemma it suffices to show that the left and right vertical maps are injective. In [Bak-computation-L-groups], Bak gave computations of $L_{i}^{x}(\mathbb{Z}G)$ for $x\in\{\mathrm{s},\mathrm{h},\mathrm{p}\}$ and $G$ a finite group whose 2-hyperelementary subgroups are abelian, which certainly holds for finite cyclic groups. See also [Bak-odd-L-groups-of-odd-order-groups, Bak-even-L-groups-of-odd-order-groups]. By [Bak-computation-L-groups, Theorem 8], the forgetful map $L^{\mathrm{h}}_{2k-1}(\mathbb{Z}C_{m})\to L^{\mathrm{p}}_{2k-1}(\mathbb{Z}C_{m})$ is injective, so the right vertical map is injective.
To prove that the left vertical map is injective we also use [Bak-computation-L-groups]. For a finite group $G$ whose 2-Sylow subgroup $G_{2}$ is normal and abelian, Bak defined $r_{2}:=\operatorname{rk}H^{1}(C_{2};\operatorname{Wh}(G))$. On [Bak-computation-L-groups, p. 386], he noted that if $G_{2}$ is cyclic, as in our case $G=C_{m}$, then $r_{2}=0$. Therefore by the sequence (3.2), as shown on [Bak-computation-L-groups, p. 390], it follows that $L_{2k}^{\mathrm{s}}(\mathbb{Z}C_{m})\to L_{2k}^{\mathrm{h}}(\mathbb{Z}C_{m})$ is injective. Hence $F\colon L_{2k}^{\mathrm{s}}(\mathbb{Z}G_{m})\to L_{2k}^{\mathrm{h}}(\mathbb{Z}G_{m})$ is injective as desired. ∎
Recall the definition of the sets $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$, $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ and $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M)$ from Definition 4.10. Also recall the definition of $A_{2k}(m)$ (Definition 6.5): $A_{2k}(m)\leq\operatorname{Aut}(C_{\infty}\times C_{m})$ denotes the subgroup of matrices $\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)$ such that $c^{k}\equiv\pm 1$ mod $m$. Note that $A_{2k}(m)$ acts on $\operatorname{Wh}(C_{\infty}\times C_{m})$ and on the subgroups $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ and $\mathcal{J}_{n}(C_{\infty}\times C_{m})$.
Proposition 7.2.
Let $m\geq 2$ and let $q_{1},\ldots,q_{k}$ be integers such that $\gcd(m,q_{j})=1$. Let $L=L_{2k-1}(m;q_{1},\ldots,q_{k})$ be a lens space of dimension n1=2k1n-1=2k-1 with $\pi_{1}(L)\cong C_{m}$. Let $G:=C_{\infty}\times C_{m}\cong\pi_{1}(S^{1}\times L)$. Then there are bijections of pointed sets $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(S^{1}\times L)\cong\mathcal{J}_{n}(G)/A_{2k}(m)$, $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(S^{1}\times L)\cong\mathcal{I}_{n}(G)/A_{2k}(m)$ and $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(S^{1}\times L)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/A_{2k}(m)$.
Proof.
@@
Firstly, since $C_{\infty}\times C_{m}$ is polycyclic, it is a good group [FQ, Freedman-book-goodgroups], so 1.1 is satisfied. Moreover, since the map $\psi$ is surjective for $G=C_{\infty}\times C_{m}$, $w\equiv 1$ and $n$ even by Proposition 7.1, the vertical maps in the diagram of Theorem 4.11 (c) are bijections if $M=S^{1}\times L$. Therefore $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(S^{1}\times L)\cong\mathcal{J}_{n}(G)/\operatorname{hAut}(S^{1}\times L)$, $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(S^{1}\times L)\cong q(\mathcal{I}_{n}(G))$, and $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(S^{1}\times L)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/\operatorname{hAut}(S^{1}\times L)$.
@@
By Theorem 6.2, every homotopy automorphism of $S^{1}\times L$ is simple, therefore the action of $\operatorname{hAut}(S^{1}\times L)$ on $\operatorname{Wh}(G)$ factors through the action of $\operatorname{Aut}(G)$ (see Definition 4.3 and Remark 4.18). By Theorem 6.6, the image of $\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(G)$ is $A_{2k}(m)$. In particular, the orbits of the action of $\operatorname{hAut}(S^{1}\times L)$ and $A_{2k}(m)$ on $\mathcal{J}_{n}(G)$, $\mathcal{I}_{n}(G)$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))$ coincide. This implies that $\mathcal{J}_{n}(G)/\operatorname{hAut}(S^{1}\times L)=\mathcal{J}_{n}(G)/A_{2k}(m)$, $q(\mathcal{I}_{n}(G))=\mathcal{I}_{n}(G)/A_{2k}(m)$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/\operatorname{hAut}(S^{1}\times L)=\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))/A_{2k}(m)$, as required. ∎
@@
In particular Proposition 7.2 implies that $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(S^{1}\times L)|$, $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(S^{1}\times L)|$ and $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(S^{1}\times L)|$ are independent of the choice of the $q_{j}$ and of $\operatorname{CAT}$, and only depend on $n$ and $m$. This proves part (a) of Theorems C, D, LABEL:, and E.
@@
From now on, we write $M^{n}_{m}$ for $S^{1}\times L$, where $L$ is any (n1)(n-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=1$ if and only if $\mathcal{J}_{n}(G)=0$.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ if and only if $|\mathcal{J}_{n}(G)|=\infty$.
@@
    If $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})$ is finite, then $\frac{|\mathcal{J}_{n}(G)|}{2m^{2}}<|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|\leq|\mathcal{J}_{n}(G)|$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|=1$ if and only if $\mathcal{I}_{n}(G)=0$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ if and only if $|\mathcal{I}_{n}(G)|=\infty$.
@@
    If $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})$ is finite, then $\frac{|\mathcal{I}_{n}(G)|}{2m^{2}}<|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|\leq|\mathcal{I}_{n}(G)|$.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|=1$ if and only if $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))=0$.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|=\infty$ if and only if $|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))|=\infty$.
@@
    If $\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})$ is finite, then $\frac{|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))|}{2m^{2}}<|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|\leq|\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G))|$.
@@
(a) By Proposition 7.2, $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=|\mathcal{J}_{n}(G)/A_{2k}(m)|$. The group $A_{2k}(m)\leq\operatorname{Aut}(G)$ acts on $\operatorname{Wh}(G)$, and hence on $\mathcal{J}_{n}(G)$, by automorphisms. So $0\in\mathcal{J}_{n}(G)$ is a fixed point of this action, and $|\mathcal{J}_{n}(G)/A_{2k}(m)|=1$ if and only if $\mathcal{J}_{n}(G)=0$.
@@
The proofs of parts (d), (e), and (f) (resp. parts (g), (h) and (i)) are entirely analogous to those of parts (a), (b), and (c) respectively, and so are omitted for brevity. ∎
We can now prove our main result.
Theorem A.
Proof.
Let $m\geq 2$ be an integer that is not square-free (e.g. $m=4$). By Theorem 5.8 and Lemma 5.12 $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|=\infty$. Hence by Lemma 7.3 (b) $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})$ is infinite. Note that $M^{n}_{m}=S^{1}\times L$ is orientable, so the same is true for every $n$-manifold homotopy equivalent to it. Therefore we can get a suitable infinite collection by choosing a representative from each element of $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})$. ∎
The following completes the proof of the remaining parts of Theorems C, D, LABEL:, and E, subject to the proofs of Theorems 11.1 and 11.2 which are postponed until Section 11.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,17,19\}$.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|\to\infty$ as $m\to\infty$ ((uniformly in $n)$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free.
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|\to\infty$ as $m\to\infty$ ((uniformly in $n)$.
@@
    There are infinitely many $m$ such that $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|=1$ for every $n$.
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|$ is finite for every $n$ and $m$.
@@
    $\limsup_{m\to\infty}\Bigl(\inf_{n}|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|\Bigr)=\infty$.